# Learning Diffeomorphism for Image Registration with Time-Continuous Networks using Semigroup Regularization

### v2

@@ Abstract
Diffeomorphic image registration (DIR) is a critical task in 3D medical image analysis, aimed at finding topology preserving deformations between pairs of images. Focusing on the solution of the flow map differential equation as the diffeomorphic deformation, recent methods use discrete timesteps along with various regularization terms to penalize the negative determinant of Jacobian and impose smoothness of the solution vector field. In this paper, we propose a novel learning-based approach for diffeomorphic 3D-image registration which finds the diffeomorphisms in the time continuum with only a single regularization term and no additional integration. As one of the fundamental properties of flow maps, we exploit the semigroup property as the only form of regularization, ensuring temporally continuous diffeomorphic flows between pairs of images. Leveraging this property, our method alleviates the need for additional regularization terms and scaling and squaring integration during both training and evaluation. To achieve time-continuous diffeomorphisms, we employ time-embedded UNets, an architecture commonly utilized in diffusion models. The proposed method reveals that ensuring diffeomorphism in a continuous time interval leads to better registration results. Experimental results on four public datasets demonstrate the superiority of our model over both learning-based and optimization-based methods. (https://github.com/mattkia/SGDIR/)
@@ 1 Introduction
Diffeomorphic image registration (DIR) is a computational technique used to align images into a common coordinate system using a deformation field, ensuring that the transformation is smooth and invertible [1, 2]. Unlike simpler forms of image registration that might allow folding or tearing, diffeomorphic methods preserve the topological properties of the image since the mapping between corresponding points is continuous, invertible, and differentiable. This property is significant in medical imaging, where anatomical structures need to be aligned accurately without grid folding to facilitate precise analysis and diagnosis.
@@
Addressing the problem of diffeomorphic image registration has led to the development of various approaches, broadly categorized into traditional, optimization-based, and learning-based methods. Traditional methods such as SyN [3], NiftyReg [4], B-Splines [5], and LDDMM [6] usually start with an assumption on the model functional form and try to maximize the similarity between the moving and the fixed images while constraining the model to satisfy some regularity properties. Even though such methods can usually generate fine detailed registrations, they suffer from relatively long computation times. Recent advancements in deep neural networks have yielded optimization-based approaches using neural networks which alleviate the restriction of having an assumption on the model functional and replace the registration model with a neural network with more expressive power [7, 8, 9, 10, 11, 12]. Such models have demonstrated superior performance and lower computation times compared to traditional methods. On the other hand, learning-based approaches [13, 14, 15, 16, 17, 18, 19, 20, 21] have gained traction due to their ability to learn implicit representations of the registration process from large datasets. These methods, still leveraging deep neural network architectures can offer potentially more robust solutions, often significantly reducing computation time compared to traditional and neural optimization-based methods.
@@
Instead of directly modeling the deformation field, most of optimization- and learning-based methods [9, 10, 19] model the velocity field controlling the dynamics of the deformation and use scaling and squaring integration scheme [22, 23] to reconstruct the deformation. Such methods typically regularize their model with various regularization terms to directly penalize the foldings in the deformation and impose smoothness on the vector field gradients and magnitudes. The advent of time-embedded architectures and time-embedded UNets [24, 25, 26] has facilitated the incorporation of the notion of time into the neural networks. Models such as DiffuseMorph [27] directly employ the Diffusion Model architecture to directly and continuously learn the deformation.
@@
In this paper, we present a novel DIR method, called SGDIR (Semi-Group DIR), which is capable of directly generating the continuous deformation without using scaling and squaring integration and multiple regularization terms. Our method is built upon a simple yet important property of flow maps, called the semi-group property [28] which governs the composition rule of flows along time. Using time-embedded UNets our model is able to retrieve the deformation at any time step instantaneously, and hence provide a smooth transition, diffeomorphic at any time step, from the moving image to the fixed image. This is unlike the DiffuseMorph which requires to start from time zero and build up the deformation up to the given time due to following diffusion models paradigm. With a novel regularization term using the semi-group property of flow maps, we ensure diffeomorphic deformations without using any other regularization terms including penalizing negative Jacobian determinants of the deformation. Furthermore, our formulation allows us to simultaneously retrieve the reverse deformation from the fixed image to the moving image.
Our contributions can be summarized as the followings:
@@
  • We introduce semi-group regularization as the only regularization term for leading the model to generate smooth and invertible deformations.
  • We provide theoretical properties and consequences of the semi-group regularization including how it leads to temporally-continuous diffeomorphic deformations.
  • We provide extensive experiments on four 3D image registration datasets including brain and lung MR scans, which demonstrate the outperformance of SGDIR with respect to recent DIR methods.
@@ 2.1 Pairwise Optimization-Based Methods
Traditional methods find the deformation field between a pair of images by minimizing an energy functional while restricting the solution space by imposing some regularity conditions on the model. SyN [6] as a successful traditional model provides a greedy technique for a symmetric diffeomorphic deformation solution by minimizing the cross-correlation between the pair of images. NiftyReg [4] as a powerful package for image registration considers different sets of linear and non-linear (such as B-Splines and Free Form Deformations (FFD)) parametric models [29] and minimizes loss functions such as Normalized Mutual Information (NMI) and Sum of Squared Differences (SSD) to maximize the alignment between the pairs of images. Traditional methods such as Demons and its variants [30, 31, 32] are built upon optical flows. LDDMM [6] is a seminal work in flow-based registration which models the registration as the geodesic paths in the images space and is proper for large deformations while preserving the topology. DARTEL [33] is also another flow-based approach which models the deformations using exponentiated Lie algebras and utilizes the scaling and squaring integration to reconstruct the deformations.
Neural optimization based approaches model the registration using neural networks. IDIR [7] and DNVF [9] are coordinate-based registration methods which utilize implicit neural representations [34]. IDIR directly learns the deformation and applies Jacobian regularization for better topology preservation and hyperelastic [35] regularization for smooth deformation, while DNVF implicitly learns the velocity field and using scaling and squaring integration reconstructs the deformation. DNVF uses Jacobian regularization along with vector field gradient and magnitude regularization. NODEO [8] which also works with voxel coordinate clouds, incorporates neural ODE networks [36] to learn the dynamics of the deformation by estimating the velocity field and finding the deformation using an ODE solver.
@@ 2.2 Learning-Based Methods
Unlike optimization-based methods which perform pairwise optimization, learning-based methods try to learn the deformation across large training datasets. The most advantageous aspect of learning-based methods is their short inference time which is orders of magnitude shorter than optimization based methods. As a fundamental work, VoxelMorph [20] directly models the deformation using a UNet and by using localized Normalized Cross Correlation metric and smoothness regularization learns the deformation fields. SYMNet [19] learns the velocity fields of forward and backward deformation and using scaling and squaring integration reconstructs the deformation. SYMNet uses Jacobian and smoothness regularizations along with minimizing the discrepancy between the forward and backward velocity fields, leading to a smoother deformation. LapIRN [37] introduces multi-resolution deformations for better generalization and avoiding local minima, and cLapIRN [16] extends LapIRN and feeds the smoothness regularization factors to the layers of the network. MS-ODE [38] incorporates neural ODEs as a refinement stage by modeling the dynamics of the parameters of a registration model.
DiffuseMorph [27] utilizes denoising diffusion models to learn temporally continuous deformations. It also incorporates Jacobian and smoothness regularization for imposing regularity on the solution. However using diffusion models could be challenging due to expensive training and maintaining visually pleasing outputs. Also, since DiffuseMorph follows the paradigm of diffusion models it requires fixed number of time samples during both training and evaluation. Furthermore, if queried at different time steps, DiffuseMorph requires to build the deformation from time step zero up to the queried time steps. Similar to DiffuseMorph we use time-embedded UNets to learn the temporally continuous deformations, but there are fundamental differences between our method and DiffuseMorph: we use a much simpler UNet without any attention modules; we are training the model with continuous time frame and we can query the model at any time step instantly. Moreover, we only use semigroup regularization introduced in this paper for learning diffeomorphic deformations.
@@ 3 Preliminaries
Deformable registration for 3D images seeks for a vector field $\phi:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ which when applied to the moving image $I_{m}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$, deforms it smoothly towards the fixed image $I_{f}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$. $\Omega\subset\mathbb{R}^{3}$ is usually a cubical region. In diffeomorphic image registration we additionally require the deformation field to be an orientation-preserving diffeomorphism, which can be mathematically translated to $\phi$ having positive determinant of Jacobian $|J_{\phi}|$ at all points. The deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ (we will use this definition going forward) can also be considered to be the time-continuous flow map solution of the following autonomous ordinary differential equation [39, 40, 6, 33, 19]:
@@
where $\phi_{0}(x)$ is short for $\phi(x,t=0)$ and $\mathbf{v}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ is a stationary velocity field governing the dynamics of the flow. In the context of DIR, $\phi_{1}(x)$ is taken to be the deformation which warps $I_{m}$ to $I_{f}$. The utility of the ODE of Eq.(3) is that its solution is guaranteed to be a diffeomorphism. It is shown that considering
@@
$\phi_{1/2^{N-1}}=\phi_{1/2^{N}}\circ\phi_{1/2^{N}}\Rightarrow\phi_{1/2^{N-2}}=\phi_{1/2^{N-1}}\circ\phi_{1/2^{N-1}}\Rightarrow\dots\Rightarrow\phi_{1}=\phi_{1/2}\circ\phi_{1/2}.$ (5)
Conventional DIR methods, taking advantage of the above properties, assume the following general loss function to address the diffeomorphic registration problem
@@
where $\mathcal{D}$ is the distribution of the images, $\mathcal{L}_{\text{sim}}(.,.)$ is a measure of similarity between the fixed image and the warped moving image, and $\mathcal{L}_{\text{reg}}(\phi)$ contains restrictive regularization terms over the deformation $\phi$. Here, $\phi[I]$ means the warping of image $I$ under the deformation field $\phi$. Starting with an estimation of the vector field $\mathbf{v}$ in Eq.(4), the deformation $\phi$ (more precisely, $\phi_{1}$) is obtained by scaling and squaring integration, and the whole model is optimized to minimize the loss Eq. (6).
On the other hand, a necessary and sufficient condition of $\phi$ as the flow map solution of the ODE of Eq.(3) is that it satisfies the semigroup property; i.e., for any time steps $t$ and $s$ it holds [28]:
@@
where $Id$ denotes identity function. Intuitively, semigroup property implies that deforming an image $I$ up to a time $t+s$ is equivalent to first deforming the image up to time $t$ to obtain $I_{t}$, and then applying the deformation again up to time $s$ given that the initial condition has changed to $I_{t}$. Also, since $\phi_{t}\circ\phi_{-t}(x)=\phi_{-t}(x)\circ\phi_{t}=\phi_{0}(x)=x$, we deduce that
@@
for all values of $x$. Given that $\phi_{t}$ is a differentiable function, Eq. (9) guaranties the bijectivity of $\phi_{t}(x)$, and hence, the diffeomorphic property of $\phi_{t}$. In the next section we’ll show how we exploit the semigroup property of Eq.(8) to circumvent the scaling and squaring integration and other regularization terms and perform the registration in a continuous time frame in contrast to former DIR methods which find the deformation at time $1$.
@@ 4 Proposed Method
Here we elaborate our method, dubbed as SGDIR. Let $I_{f},I_{m}:\Omega\longrightarrow\mathbb{R}$ be the fixed and moving images, respectively, sampled from the images distribution $\mathcal{D}$. $\Omega\subset\mathbb{R}^{3}$ is a $D\times H\times W$ cubical region. We are looking for a time-continuous deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ to deform the moving image towards the fixed image, and vice versa. We assume that $\phi$ is the flow map solution of the ODE (3), for some unknown stationary vector field $\mathbf{v}$, and therefore, it must satisfy the semigroup property of Eq.(8). If we warp $I_{m}$ up to time $t$ using the deformation $\phi_{t}$, and warp $I_{f}$ up to time $1-t$ using the inverse deformation $\phi^{-1}_{1-t}$, we must reach to a same point due to the continuity of the trajectory of $\phi$. Based on Eq.(9) on the inversion of $\phi$, we can write
Figure 1: The general schematic of the SGDIR training. As illustrated in the figure, there are no additional integration or extra loss functions applied which makes the whole procedure simple. During inference, we can query the model for any desired $t$, e.g., $t=1$ for the final deformation.
@@
Eq.(10) allows us to define time-continuous similarity loss. For measuring the similarity, we use the localized Normalized Cross Correlation (NCC) between the images [20, 8] (see Appendix 7.1).
In this work we use a local window of size $11$ in NCC. Using Eq.(10), we introduce a time-continuous similarity loss:
@@
for all $t\in[-1,1]$, where $\mathbf{F}(x,t;I_{f},I_{m},\theta)$ is a TE-UNet with learnable parameters $\theta$ which also receives a pair of fixed and moving images. The reason behind the specific formulation of Eq.(12) is that we can make sure that at $t=0$ we have $\phi_{0}(x)=x$, hence satisfying the identity initial condition of the ODE (3). In order for $\phi$ to be a valid flow map, it needs to satisfy the semigroup property of Eq. (8). In practice, forcing the model to perfectly satisfy the semigroup property could be cumbersome and expensive, since we need to sample $t$ and $s$ independently and impose Eq. (11) twice, which requires more forward calls from the network. To tackle this issue we propose the following regularization,
$\mathcal{L}^{t}_{\text{reg}}=\|\phi_{2t-1}-\phi_{t}\circ\phi_{t-1}\|_{2}+\|\phi_{2t-1}-\phi_{t-1}\circ\phi_{t}\|_{2},$ (13)
for all $t\in[0,1]$, and we show that it suffices to render $\phi$ invertible and consequently a diffeomorphism at all time steps $t\in[-1,1]$. In Eq. (13), .$\|.\|_{2}$ denotes the $L_{2}$-norm, and forces the model to satisfy the semigroup property for $t$ and $s=t-1$. The second term in Eq. (13) is added only to impose symmetry and provide diffeomorphism in $[-1,0]$ interval. Similar to all DIR methods we implement the composition operation by trilinear interpolation of grids over each other [19, 9, 37, 16]. From Eq. (13), one could instantly say that a model that minimizes Eq. (13), will satisfy the following rules
@@
The deformation $\phi(x,t)$ optimally learned by minimizing Eq. (11) and Eq. (13), is a diffeomorphism at any time $t\in[-1,1]$.
@@
Finally, we can propose the following loss function which is in concordance with Eq. (6):
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D},t\sim Uni(0,1)}\left[\mathcal{L}^{t}_{\text{sim}}+\lambda\mathcal{L}^{t}_{\text{reg}}\right],$ (15)
where $Uni(0,1)$ is the uniform distribution on $[0,1]$ and $\mathcal{L}^{t}_{\text{sim}}$ and $\mathcal{L}^{t}_{\text{reg}}$ are defined as in Eq. (11) and Eq. (13), respectively. $\lambda$ is the regularization factor controlling the level of diffeomorphism. In the training phase, we randomly sample a pair of images to be registered along with a time step uniformly taken from the interval $[0,1]$ to minimize Eq. (15). The trained model is capable of warping either $I_{f}$ or $I_{m}$ towards each other at any time step. Thus, the model is more versatile compared to the most previous works that provide the deformation in only one direction.
It is also worth mentioning that some prior works [41, 42] enforce forward-backward deformation consistency as a regularization term by constructing a time-independent forward deformation only at the trajectory endpoint ($\phi_{1}$), while explicitly modeling the backward deformation ($\phi_{1}^{-1}$). In contrast, SGDIR allows evaluating the deformation field $\phi_{t}$ at arbitrary time steps, without separately modeling the inverse. Consistency between forward and backward deformations, as we prove, is obtained as a natural result of semigroup regularization, rather than through direct imposition. LDDMM [6] models a time-dependent deformation using discrete time steps as opposed to our continuous time steps and uses forward-backward calculations to compute the velocity field explicitly. In contrast, we do not compute any velocity field explicitly.

### v3

@@ Abstract
Diffeomorphic image registration (DIR) is a fundamental task in 3D medical image analysis that seeks topology-preserving deformations between image pairs. To ensure diffeomorphism, a common approach is to model the deformation field as the flow map solution of a differential equation, which is solved using efficient schemes such as scaling and squaring along with multiple smoothness regularization terms. In this paper, we propose a novel learning-based approach for diffeomorphic 3D image registration that models diffeomorphisms in a continuous-time framework using only a single regularization term, without requiring additional integration. We exploit the semigroup property—a fundamental characteristic of flow maps—as the sole form of regularization, ensuring temporally continuous diffeomorphic flows between image pairs. Leveraging this property, we prove that our formulation directly learns the flow map solution of an ODE, ensuring continuous inverse and cycle consistencies without explicit enforcement, while eliminating additional integration schemes and regularization terms. To achieve time-continuous diffeomorphisms, we employ time-embedded UNets, an architecture commonly used in diffusion models. Our results demonstrate that modeling diffeomorphism continuously in time improves registration performance. Experimental results on four public datasets demonstrate the superiority of our model over state-of-the-art diffeomorphic methods. Additionally, comparison to several recent non-diffeomorphic deformable image registration methods shows that our method achieves competitive Dice scores while significantly improving topology preservation. (https://github.com/mattkia/SGDIR/)
@@ 1 Introduction
Diffeomorphic image registration (DIR) is a computational technique used to align images into a common coordinate system using a deformation field, ensuring that the transformation is smooth and invertible [1, 2, 3]. Unlike simpler forms of image registration that might allow folding or tearing, diffeomorphic methods preserve the topological properties of the image since the mapping between corresponding points is continuous, invertible, and differentiable. This property is crucial in medical imaging, where anatomical structures need to be aligned accurately without grid folding to facilitate precise analysis and diagnosis.
@@
Addressing the problem of diffeomorphic image registration has led to the development of various approaches, broadly categorized into traditional, instance optimization-based, and learning-based methods. Traditional methods such as SyN [4], NiftyReg [5], B-Splines [6], and LDDMM [7] usually start with an assumption on the model functional form and try to maximize the similarity between the moving and the fixed images while constraining the model to satisfy some regularity properties. Even though such methods can usually generate fine-detailed registrations, they suffer from relatively long computation times. Recent advances in deep neural networks have yielded instance optimization-based approaches using neural networks which alleviate the restriction of having an assumption on the model functional and replace the registration model with a neural network with more expressive power [8, 9, 10, 11, 12, 13]. Such models have demonstrated superior performance and lower computation times compared to traditional methods. On the other hand, learning-based approaches [14, 15, 16, 17, 18, 19, 20, 21] have gained traction due to their ability to learn implicit representations of the registration process from large datasets. These methods, still leveraging deep neural network architectures can offer potentially more robust solutions, often significantly reducing inference time compared to traditional and neural optimization-based methods.
@@
Instead of directly modeling the deformation field, most of optimization- and learning-based methods [10, 11, 20] model the velocity field controlling the dynamics of the deformation and use scaling and squaring integration scheme to reconstruct the deformation [22, 23]. Such methods typically regularize their model with various regularization terms to directly penalize the foldings in the deformation and impose smoothness on the vector field gradients and magnitudes. The advent of time-embedded architectures and time-embedded UNets [24, 25, 26] has facilitated the incorporation of the notion of time into the neural networks. Models such as DiffuseMorph and DiffuseReg [27, 28] directly employ the Diffusion Model architecture to directly and continuously learn the deformation.
It is also known that the inverse consistency of a registration model plays a critical point in the behavior of the registration method especially for the medical image registration task [29]. To this end, various methods have incorporated cycle and inverse consistencies either into the model’s architecture or in their training procedure [29, 30, 31, 32].
@@
In this paper, we present a novel DIR method, called SGDIR (Semi-Group DIR), which is capable of directly generating the continuous deformation without using scaling and squaring integration and multiple regularization terms. Our method is built upon a simple yet important property of flow maps, called the semigroup property [33] which governs the composition rule of flows along time. Using time-embedded UNets our model is able to retrieve the deformation at any time step instantaneously, and hence provides a smooth transition, diffeomorphic at any time step, from the moving image to the fixed image, and vice versa. We will prove that our particular semigroup regularization ensures that the model learns a deformation field that is indeed the solution to an ODE, hence, implicitly satisfying the diffeomrophic property and cycle and inverse consistencies, without using any other regularization terms including penalizing negative Jacobian determinants of the deformation. Our contributions can be summarized as the followings:
@@
  • We introduce a simple semigroup regularization as the only regularization term and we mathematically prove that our regularization term leads the model towards learning the solution of the deformation ODE, hence, implicitly satisfying the diffeomorphic property.
@@
  • We provide extensive experiments on four 3D image registration datasets on brain MR scans, which demonstrate the superior accuracy of SGDIR compared to recent DIR methods. When compared with recent non-diffeomorphic deformable methods, SGDIR show significant improvement in topology preservation without sacrificing accuracy.
@@ 2.1 Pairwise Optimization-Based Methods
Traditional methods find the deformation field between a pair of images by minimizing an energy functional while restricting the solution space by imposing some regularity conditions on the model. SyN [7] as a successful traditional model provides a greedy technique for a symmetric diffeomorphic deformation solution by minimizing the cross-correlation between the pair of images. NiftyReg [5] as a powerful package for image registration considers different sets of linear and non-linear parametric models [34] and minimizes loss functions such as Normalized Mutual Information (NMI) and Sum of Squared Differences (SSD) to maximize the alignment between the pairs of images. Traditional methods such as Demons and its variants [35, 36, 37] are built upon optical flows. LDDMM [7] is a seminal work in flow-based registration that models the registration as the geodesic paths in the images space and is proper for large deformations while preserving the topology. DARTEL [38] is also another flow-based approach which models the deformations using exponentiated Lie algebras and utilizes scaling and squaring integration to reconstruct the deformations.
Neural optimization based approaches model registration using neural networks. IDIR [8], DNVF [10], and NePhi [39] are coordinate-based registration methods which utilize implicit neural representations [40]. IDIR directly learns the deformation and applies Jacobian regularization for better topology preservation and hyperelastic regularization for smooth deformation [41], while DNVF implicitly learns the velocity field and using scaling and squaring integration, reconstructs the deformation. DNVF uses Jacobian regularization along with vector field gradient and magnitude regularization. NODEO [9] which also works with voxel coordinates, incorporates neural ODEs [42] to learn the dynamics of the deformation by estimating the velocity field and finding deformation using an ODE solver.
@@ 2.2 Learning-Based Methods
Unlike optimization-based methods which perform pairwise optimization, learning-based methods try to learn the deformation across large training datasets. The most advantageous aspect of learning-based methods is their short inference time which is orders of magnitude shorter than that of optimization based methods. As a pioneering work, VoxelMorph [21] directly models the deformation using a UNet and by using localized Normalized Cross Correlation metric and smoothness regularization learns the deformation fields. SYMNet [20] learns the velocity fields of forward and backward deformation and using scaling and squaring integration reconstructs the deformation. SYMNet uses Jacobian and smoothness regularizations along with minimizing the discrepancy between the forward and backward velocity fields, leading to a smoother deformation. LapIRN [43] and PULPo [44] employ multi-resolution deformations for better generalization and avoiding local minima. cLapIRN [17] extends LapIRN and feeds the smoothness regularization factors to the layers of the network. MS-ODE [45] incorporates neural ODEs as a refinement stage by modeling the dynamics of the parameters of a registration model. TransMorph [46] introduces a general deformable image registration method by utilizing Vision Transformers, and further they customize their model for diffeomorphic registration, known as TransMorph-diff. H-ViT also uses the vision transformers along with cross-attention mechanism to produce more accurate deformations [47].
DiffuseMorph [27] utilizes denoising diffusion models to learn temporally continuous deformations. It also incorporates Jacobian and smoothness regularizations for imposing regularity on the solution. Instead of focusing on the images, DiffuseReg denoises the deformation field and follows the diffusion models strategy [28]. However, using diffusion models could be challenging due to expensive training, maintaining visually pleasing outputs, and slow generation. We also use time-embedded UNets to learn the temporally continuous deformations. Unlike the diffusion model paradigm, we are able to output the deformation instantly at any timestep instead of building it from time step zero. Our simple design endows us with the ability to generate the deformation fields significantly faster.
@@ 3 Preliminaries
Figure 1: Schematic of SGDIR. As illustrated in the figure, there are no additional integration or extra loss functions to make the whole training procedure simple. During inference, we can query the model for any desired $t$, e.g., $t=1$ for the final deformation.
Deformable registration for 3D images seeks for a vector field $\phi:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ that when applied to the moving image $I_{m}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$, deforms it smoothly towards the fixed image $I_{f}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$. The domain $\Omega\subset\mathbb{R}^{3}$ is usually a cubical region. In diffeomorphic image registration, we additionally require the deformation field to be an orientation-preserving diffeomorphism, which can be mathematically translated to $\phi$ having positive determinant of Jacobian $|J_{\phi}|$ at all points. The deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ (we will use this definition going forward) can also be considered to be the time-continuous flow map solution of the following autonomous ordinary differential equation [48, 49, 7, 38, 20]:
@@
where $\phi_{0}(x)$ is short for $\phi(x,t=0)$ and $\mathbf{v}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ is a stationary velocity field that governs the dynamics of the flow. In the context of DIR, $\phi_{1}(x)$ is taken to be the deformation that warps $I_{m}$ to $I_{f}$. The utility of the ODE (3) is that its solution is guaranteed to be a diffeomorphism. It is shown that considering
@@
$\phi_{\frac{1}{2^{N-1}}}=\phi_{\frac{1}{2^{N}}}\circ\phi_{\frac{1}{2^{N}}}\Rightarrow\phi_{\frac{1}{2^{N-2}}}=\phi_{\frac{1}{2^{N-1}}}\circ\phi_{\frac{1}{2^{N-1}}}\Rightarrow\dots\Rightarrow\phi_{1}=\phi_{1/2}\circ\phi_{1/2}.$ (5)
Conventional DIR methods, taking advantage of the above properties, assume the following general loss function to address the diffeomorphic registration problem.
@@
where $\mathcal{D}$ is the distribution of the training images, $\mathcal{L}_{\text{sim}}(.,.)$ is a measure of similarity between the fixed image and the warped moving image, and $\mathcal{L}_{\text{reg}}(\phi)$ contains restrictive regularization terms over the deformation $\phi$. Here, $\phi[I]$ means the warping of the image $I$ under the deformation field $\phi$. Starting with an estimation of the vector field $\mathbf{v}$ in Eq. 4, the deformation $\phi$ (more precisely, $\phi_{1}$) is obtained by scaling and squaring integration, and the whole model is optimized to minimize the loss Eq. 6.
On the other hand, a necessary and sufficient condition of $\phi$ as the flow map solution of the ODE (3) is that it satisfies the semigroup property; i.e., for any time steps $t$ and $s$ it holds [33]:
@@
where $Id$ denotes identity function. Intuitively, semigroup property implies that deforming an image $I$ up to a time $t+s$ is equivalent to first deforming the image up to time $t$ to obtain $I_{t}$, and then applying the deformation again up to time $s$ given that the initial condition has changed to $I_{t}$. Also, since $\phi_{t}\circ\phi_{-t}(x)=\phi_{-t}\circ\phi_{t}(x)=\phi_{0}(x)=x$, we deduce that
@@
for all values of $x$. Given that $\phi_{t}$ is a differentiable function, Eq. 9 guaranties the bijectivity of $\phi_{t}(x)$, and hence, the diffeomorphic property of $\phi_{t}$. In the next section we’ll show how we exploit the semigroup property of Eq. 8 to circumvent the scaling and squaring integration and other regularization terms and perform the registration in a continuous time frame in contrast to former DIR methods which find the deformation at time $1$.
@@ 4 Proposed Method
Here we elaborate our method, dubbed as SGDIR. Let $I_{f},I_{m}:\Omega\longrightarrow\mathbb{R}$ be the fixed and moving images, respectively, sampled from the image distribution $\mathcal{D}$. The domain $\Omega\subset\mathbb{R}^{3}$ is a $D\times H\times W$ cubical region. We are looking for a time-continuous deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ to deform the moving image towards the fixed image, and vice versa. We assume that $\phi$ is the flow map solution of the ODE (3), for some unknown stationary vector field $\mathbf{v}$, and therefore, it must satisfy the semigroup property of Eq. 8. If we warp $I_{m}$ up to time $t$ using the deformation $\phi_{t}$, and warp $I_{f}$ up to time $1-t$ using the inverse deformation $\phi^{-1}_{1-t}$, we must reach to a same point due to the continuity of the trajectory of $\phi$. Based on Eq. 9, we can write
@@
Eq. 10 allows us to define a time-continuous similarity loss. For measuring the similarity, we use the localized Normalized Cross Correlation (NCC) between the images [21, 9] (see Appendix 7.3). In this work we use a local window of size $11$ for NCC. Using Eq. 10, we introduce a time-continuous similarity loss:
@@
for all $t\in[-1,1]$, where $\mathbf{F}(x,t;I_{f},I_{m},\theta)$ is a TE-UNet with learnable parameters $\theta$ which also receives a pair of fixed and moving images. The reason behind the specific formulation of Eq. 12 is that we can make sure that at $t=0$ we have $\phi_{0}(x)=x$, hence satisfying the identity initial condition of the ODE (3). In order for $\phi$ to be a valid flow map, it needs to satisfy the semigroup property of Eq. 8. In practice, forcing the model to perfectly satisfy the semigroup property could be cumbersome and expensive, since we need to sample $t$ and $s$ independently and impose Eq. 11 twice, which requires more forward calls from the network. To tackle this issue we propose the following simplified semigroup regularization,
$\mathcal{L}^{t}_{\text{semigroup}}=\|\phi_{2t-1}-\phi_{t}\circ\phi_{t-1}\|_{2}+\|\phi_{2t-1}-\phi_{t-1}\circ\phi_{t}\|_{2},$ (13)
for all $t\in[0,1]$, and we show that it suffices to render $\phi$ a diffeomorphism at all time steps $t\in[-1,1]$. In Eq. 13, .$\|.\|_{2}$ denotes the $L_{2}$-norm, and forces the model to satisfy the semigroup property for $t$ and $s=t-1$. The second term in Eq. 13 is added to impose symmetry and is essential to hold the semi-group property. Similar to all DIR methods we implement the composition operation by trilinear interpolation of grids over each other [20, 10, 43, 17]. From Eq. 13, we can say that a model that minimizes Eq. 13, will satisfy the following rules
@@
The deformation $\phi(x,t)$ that satisfies the composition rule of Eq. 14 and $\phi(x,0)=x$ is an exponential map, or equivalently, it is a one-parameter diffeomorphism solving an autonomous ODE (3)(see [50]).
@@
Finally, we can propose the following loss function which is in concordance with Eq. 6:
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D},t\sim Uni(0,1)}\left[\mathcal{L}^{t}_{\text{sim}}+\lambda\mathcal{L}^{t}_{\text{semigroup}}\right],$ (15)
where $Uni(0,1)$ is the uniform distribution on $[0,1]$ and $\mathcal{L}^{t}_{\text{sim}}$ and $\mathcal{L}^{t}_{\text{semigroup}}$ are defined as in Eq. 11 and Eq. 13, respectively. $\lambda$ is the regularization factor controlling the level of diffeomorphism. In the training phase, we randomly sample a pair of images to be registered along with a time step uniformly taken from the interval $[0,1]$ to minimize Eq. 15. The trained model is capable of warping either $I_{f}$ or $I_{m}$ towards the other at any time step. Thus, the model is more versatile compared to the most previous works that provide the deformation in only one direction.
Moreover, note that since the implementation of $\phi_{t}$ as explained above has an inverse for any $t\in[-1,1]$ the network actually learns a continuous group structure of deformations over the interval $[-1,1]$.
It is also worth mentioning that some prior works [31, 32, 30] enforce forward-backward deformation consistency as a regularization term by constructing a time-independent forward deformation only at the trajectory endpoint ($\phi_{1}$), while explicitly modeling the backward deformation ($\phi_{1}^{-1}$). In contrast, SGDIR allows evaluating the deformation field $\phi_{t}$ at arbitrary time steps, without separately modeling the inverse. Consistency between forward and backward deformations, as we prove, is obtained as a natural result of semigroup regularization,
Figure 2: Qualitative comparison of SGDIR with other methods on the OASIS dataset. (visually looks better in color and zoomed in.)